If X^2+ bX +c and X^2+ dX + e have a common factor ( X – p ) show that p =e-c/b-d


1
Expert's answer
2021-09-15T03:16:50-0400

If "x^2+ bx +c" has a factor "(x-p)," then


"p^2+bp+c=0"

If "x^2+ dx +e" has a factor "(x-p)," then


"p^2+dp+e=0"

Hence


"p^2+bp+c=p^2+dp+e"

"bp+c=dp+e"

"(b-d)p=e-c"

"p=\dfrac{e-c}{b-d}, b\not=d"

If "b=d," then "c=e" and the quadratic polynomials are the same for "p\in\R."